# Nonstrange and strange quark matter at finite temperature within modified NJL model and protoquark stars

### v1

@@ V.1 Strange proto-quark star properties
In general, increasing the value of $\alpha$ enhances the repulsive vector interactions, leading to a stiffer EOS, which can be seen clearly in the comparison between $\alpha=0.5$ and $\alpha=0.8$. Previous studies have also shown that strong vector interactions are necessary to support massive compact stars 2005PhR…407..205B . For large values of $\alpha$, such as $\alpha=0.8$, the chiral phase transition becomes a crossover. This smooth transition typically results in a stiffer EOS compared to the case of a first-order transition. For the case of $\alpha=0.8,Y_{l}=0.1,\ B=133\;\text{MeV}$, increasing the temperature from $T=10\;\text{MeV}$ to $40\;\text{MeV}$ slightly softens the EOS, leading to a modest reduction in the maximum mass $M_{\mathrm{TOV}}$, while increasing the stellar radius at the same mass. Because the increase in energy exceeds the thermal contribution to pressure. A similar trend is also reported in Ref. Pal & Chaudhuri (2024), where, within the quasiparticle model incorporating vector interactions, the maximum mass decreases with increasing temperature in the isothermal scenario. When the total lepton fraction $Y_{l}$ increases from 0.1 to 0.4 (with $\alpha=0.8$, $T=0.01~\mathrm{MeV}$, and $B=100~\mathrm{MeV}$), the EOS exhibits a more complex behavior due to the intersection of the two curves. At higher densities, the EOS with a larger lepton fraction becomes stiffer, primarily due to the suppression of the strange quark population.
At $\alpha=0.8$, $Y_{l}=0.1$, and $T=10\;\text{MeV}$, a representative vacuum pressure of $B=110,\mathrm{MeV}$ yields a quark star model consistent with observational constraints from both the NICER mission 2019ApJ…887L..24M ; 2019ApJ…887L..21R ; 2021ApJ…918L..28M ; 2021ApJ…918L..27R and the compact object located in the supernova remnant HESS J1731–347 Doroshenko et al. (2022). Further reducing the vacuum pressure to $B=100\;\text{MeV}$ results in an even stiffer EOS, leading to a particularly intriguing outcome—the appearance of a twin-star configuration Alford et al. (2013); Alvarez-Castillo & Blaschke (2017); Christian et al. (2018); Montaña et al. (2019), i.e., two compact stars with nearly the same gravitational mass but substantially different radii, located on two separate stable branches of the mass–radius relation. This phenomenon can be attributed to the emergence of strange quarks at typical baryon number densities. As shown in Fig. 4, the appearance of strange quarks initially softens the EOS slightly, followed by a stiffening at higher densities. As discussed in Refs. Alford et al. (2013); Alvarez-Castillo & Blaschke (2017); Christian et al. (2018); Montaña et al. (2019), a sharp first-order phase transition inside compact stars can lead to twin-star configurations. In our study, we find that such configurations can also arise in some parameter sets even when the chiral transition is a smooth crossover. It should be noted that the gravitational wave constraints from GW170817 were derived from the late inspiral phase of the merger, under the assumption of zero temperature 2017PhRvL.119p1101A ; 2018PhRvL.121p1101A . Therefore, while we include these constraints for reference, they are not directly applicable to hot compact stars. The precise temperature of neutron stars during the inspiral phase remains an open question Laskos-Patkos et al. (2022). In the near future, we will investigate the influence of finite-temperature effects on tidal deformability and the I-Love-Q relations within the modified NJL models.
V.2 Nonstrange proto-quark star properties
Figure 5 shows the EOS of nonstrange quark matter derived from our modified NJL model, together with the corresponding mass–radius relations for several representative parameter sets. For both $\alpha=0.8$ and $\alpha=0.9$, the chiral phase transition is smooth, yielding continuous EOSs. To ensure that finite-temperature nonstrange quark matter is self-bound—i.e. capable of forming a quark star—we add a vacuum pressure term $B$. At $\alpha=0.8,\ T=10\;\text{MeV},\ B=80\;\text{MeV}$, raising the lepton fraction from $Y_{l}=0.1$ to $Y_{l}=0.4$ stiffens the EOS significantly, so that the resulting maximum masses are compatible with NICER constraints. Increasing the exchange channel strength from $\alpha=0.8$ to $\alpha=0.9$ likewise produces a marked stiffening of the EOS. Reducing the vacuum pressure from $B=80\;\text{MeV}$ to $B=60\;\text{MeV}$ not only stiffen the EOS but also affects the low-density regime, leading to larger stellar radii than in the $B=80\;\text{MeV}$ case.
Different with the case of (2+1)-flavor quark matter, increasing the temperature of nonstrange quark matter tends to stiffen the EOS, as the rise in pressure outweighs the thermal contribution to the energy density. For example, with $\alpha=0.8,Y_{l}=0.4,\ B=80\;\text{MeV}$, increase the temperature from $10\;\text{MeV}$ to $40\;\text{MeV}$ raises the maximum mass $M_{\rm TOV}$ from $1.93\,M_{\odot}$ to $2.05\,M_{\odot}$. The influence of temperature on nonstrange quark matter is analogous to that observed in neutron matter, leading to a stiffer EOS and an increased maximum mass at higher temperatures Koliogiannis & Moustakidis (2021); Wu et al. (2025). To remain consistent with current astronomical observations, a sufficiently strong vector interaction and a relatively small vacuum pressure are required. No parameter space is found in which proto-nonstrange quark stars can form a twin star configuration.
@@ VI Summary
In this work, we investigate protoquark stars by combining the EOS of quark matter at finite temperature with a modified NJL model that incorporates exchange interactions, weighted by a parameter $\alpha$, for both two- and three-flavor cases. Unlike previous studies that introduce vector interactions manually, our approach uses the Fierz transformation to include scalar and vector contributions from the exchange channels self-consistently. Within this framework, we find that increasing the exchange interaction strength $\alpha$ modifies the chiral phase transition from a first-order transition to a crossover. Treating the vacuum pressure as a phenomenological parameter, we find that a first-order chiral phase transition within the NJL model necessitates a relatively large vacuum pressure to ensure vanishing pressure at the stellar surface, thereby yielding a softer EOS. An increase in the lepton fraction $Y_{l}$ stiffens the EOS for both strange and nonstrange quark matter at finite temperature. Temperature effects differ between strange and nonstrange quark matter: high temperatures tend to soften the EOS for strange quark matter, while relatively stiffening it for nonstrange quark matter. Notably, for protoquark stars composed of strange quark matter, we identify a region in parameter space that allows for the existence of twin star configurations.
Looking ahead, a more realistic treatment of protoquark stars should incorporate isotropic conditions and magnetic field effects Chu et al. (2020, 2022). Beyond the normal quark phase, the potential role of a color superconducting phase and pion condensation phase at finite temperature is also particularly intriguing—especially in scenarios where high baryon densities are realized in the early Universe due to a large lepton number Vovchenko et al. (2021); Ferreira et al. (2025). Within the same finite-temperature NJL framework, it would be highly interesting to investigate the influence of diquark condensates on the evolution of protoquark stars. The corresponding results would be crucial for understanding the cooling mechanism of the compact stars, and could shed light on the behavior of dense matter and its broader cosmological implications.
@@ Acknowledgements.
We thank Chen Zhang, Pengcheng Chu, and Bingjun Zuo for their valuable comments and useful discussions. We also thank the PKU pulsar group for the helpful discussions. This work is supported by the National SKA Program of China (2020SKA0120100), the National Natural Science Foundation of China (Nos. 12003047, 12133003), the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDB0550300). Wen-Li Yuan is supported by the Special Funds of the National Natural Science Foundation of China (Grant No. 12447171) and the China Postdoctoral Science Foundation (Grant. No. 2025M773418). Professor Nobutoshi Yasutake is supported by the JSPS KAKENHI (Grant No. 24K07054).

### v2

@@ V.1 Strange stellar quark matter and stars
In general, increasing the value of $\alpha$ enhances the repulsive vector interactions, leading to a stiffer EOS significantly, as clearly illustrated by the comparison of $\alpha=0.5$ and $\alpha=0.8$ in the behavior of the sound velocity. Strong vector interactions are necessary to support massive compact stars 2005PhR…407..205B . For large values of $\alpha$, such as $\alpha=0.8$, the chiral phase transition becomes a crossover. This smooth transition typically results in a stiffer EOS compared to the case of a first-order phase transition. For the case of $\alpha=0.8,Y_{l}=0.1,\ B^{1/4}=133\;\text{MeV}$, increasing the temperature from $T=10\;\text{MeV}$ to $40\;\text{MeV}$ softens the EOS, which can be seen from the behavior of $c_{s}^{2}$, leading to a modest reduction in the maximum mass $M_{\mathrm{TOV}}$ in Fig. 5. This reduction arises because the thermal contribution to the energy density exceeds that to the pressure. Ref. Pal & Chaudhuri (2024) has also found that the maximum mass decreases with increasing temperature in the isothermal scenario within the quasiparticle model incorporating vector interactions. When the total lepton fraction $Y_{l}$ increases from 0.1 to 0.4 (with $\alpha=0.8$, $T=10~\mathrm{MeV}$, and $B^{1/4}=95~\mathrm{MeV}$), the EOS with a larger lepton fraction becomes slightly softer, particularly at higher energy densities. At $\alpha=0.8$, $Y_{l}=0.1$, and $T=10\;\text{MeV}$, a representative vacuum pressure of $B^{1/4}=95~\;\text{MeV}$ yields a maximum mass of $1.93\,M_{\odot}$ for the strange protoquark stars. Further increasing the vacuum pressure to $B^{1/4}=133\;\text{MeV}$ results in a much softer EOS. The speed of sound directly characterizes the stiffness of the EOS. It exhibits a rapid increase at low densities, then saturates at $c_{s}^{2}\sim 0.5$, when the chiral transition undergoes a crossover at higher densities. The green dots indicate the first-order phase transition scenario, which exhibits a soft EOS with $c_{s}^{2}\sim 0.3$ across the entire density range.
As discussed in Sec. III, thermodynamic consistency requires that the minimum of the free energy per baryon, $f/\rho_{B}$, occur at zero pressure. The behavior of $f/\rho_{B}$ as a function of $P$ clearly demonstrates thermodynamic consistency within the present model and also provides insight into the nature of the phase transition. For $P>0$, the value of $f/\rho_{B}$ increases with increasing chemical potential $\mu$ (or $\rho_{B}/\rho_{0}$ and $P$). In the case of a first-order phase transition, typically realized for parameters $\alpha=0.8$, $Y_{l}=0.1$, $T=40$ MeV, and $B^{1/4}=133$ MeV, $f/\rho_{B}$ exhibits a discontinuity, whereas a smooth crossover leads to a continuous behavior. As $B^{1/4}$ decreases, the baryon number density $\rho_{B}/\rho_{0}$ corresponding to the minimum of $f/\rho_{B}$ also decreases. However, this density must be at least larger than the nuclear saturation density $\rho_{0}$ to remain consistent with terrestrial nuclear experiments, thereby providing an additional constraint on the vacuum pressure $B$. In particular, for strange quark matter with $\alpha=0.8$, $Y_{l}=0.4$, and $T=10$ MeV, a minimum value of $B^{1/4}_{\rm min}=94$ MeV is required to ensure that the baryon number density at the stellar surface satisfies $\rho_{B}>\rho_{0}$ when $P=0$. To the best of our knowledge, this physically motivated constraint on the phenomenological vacuum pressure $B$ has not been explicitly emphasized in previous studies.
It should be noted that the mass constraints derived from NICER observations 2019ApJ…887L..24M ; 2019ApJ…887L..21R ; 2021ApJ…918L..28M ; 2021ApJ…918L..27R can only be applied to cold compact stars, meanwhile the gravitational-wave constraints from GW170817 were obtained from the late inspiral phase of the merger under the assumption of zero temperature  2017PhRvL.119p1101A ; 2018PhRvL.121p1101A . Therefore, they are not directly applicable to hot compact stars. The precise temperature of neutron stars during the inspiral phase remains an open question Laskos-Patkos et al. (2022). In future work, investigating the influence of finite-temperature effects on tidal deformability and the I-Love-Q relations within the framework of the modified NJL model will further improve our understanding of dense matter and the associated astrophysical phenomena.
V.2 Nonstrange stellar quark matter and stars
The right panel of Fig. 4 shows the EOS of nonstrange quark matter derived from our modified NJL model, together with the sound velocity and the free energy per baryon for several representative parameter sets. For both $\alpha=0.8$ and $\alpha=0.9$, the chiral phase transition is smooth, yielding continuous EOSs. To ensure that finite-temperature nonstrange quark matter is self-bound, i.e. capable of forming a quark star, we add a vacuum pressure term $B$. The minima of $f/\rho_{B}$ all occur at zero pressure. To ensure that the baryon number density at the surface of the quark star exceeds $\rho_{0}$, the minimum vacuum pressure is $B^{1/4}{\rm min}=80$ MeV for $\alpha=0.8$, $Y_{l}=0.1$, and $T=10$ MeV. Increasing the strength of the exchange interaction to $\alpha=0.9$ raises the minimum vacuum pressure to $B^{1/4}{\rm min}=90$ MeV. The speed of sound quickly rises to approximately $c_{s}^{2}=0.4$ at an energy density of around $500$ MeV/fm3, with only minor changes to $c_{s}^{2}=0.46$ and $c_{s}^{2}=0.51$ for values of $\alpha$ between 0.8 and 0.9, respectively, at higher energy densities.
At $\alpha=0.8,\ T=10\;\text{MeV},\ B^{1/4}=90\;\text{MeV}$, raising the lepton fraction from $Y_{l}=0.1$ to $Y_{l}=0.4$ softens the EOS at large densities, resulting in a relatively smaller maximum mass, as shown in the right panel of Fig. 5. Increasing the exchange channel strength from $\alpha=0.8$ to $\alpha=0.9$ produces a marked stiffening of the EOS, leading to $M_{\rm TOV}=2.07$ for the maximum mass. Increasing the vacuum pressure from $B^{1/4}=90\;\text{MeV}$ to $B^{1/4}=110\;\text{MeV}$ results in a lower maximum star mass. Adding large thermal effects to the nonstrange quark matter system slightly softens the EOS, which induces a lower $M_{\rm TOV}$ but larger radius. Compared with strange quark matter, the temperature effect on the stellar radius is somewhat more pronounced in the nonstrange case.
@@ VI Summary
In this work, we investigate protoquark stars by combining the EOS of quark matter at finite temperature with a modified NJL model that incorporates exchange interactions, weighted by a parameter $\alpha$, for both two- and three-flavor cases. Unlike previous studies that introduce vector interactions manually, our approach uses the Fierz transformation to include scalar and vector contributions from the exchange channels self-consistently. Within this framework, we find that increasing the exchange interaction strength $\alpha$ modifies the chiral phase transition from a first-order transition to a crossover. Treating the vacuum pressure as a phenomenological parameter, we find that a first-order chiral phase transition within the NJL model necessitates a relatively large vacuum pressure to ensure vanishing pressure at the stellar surface, thereby yielding a softer EOS. The free energy per baryon, as well as the value of its minimum, decreases with a smaller vacuum pressure. Combining this with the requirement that the baryon number density at the stellar surface must be at least larger than $n_{0}$ provides a lower bound on the vacuum pressure $B$. In the isothermal scenario, the equation of state is almost insensitive to variations in temperature and lepton fraction.
Looking ahead, a more realistic treatment of protoquark stars should incorporate isotropic conditions and magnetic field effects Chu et al. (2020, 2022). Beyond the normal quark phase, the potential role of a color superconducting phase and pion condensation phase at finite temperature is also particularly intriguing, especially in scenarios where high baryon densities are realized in the early Universe due to a large lepton number Vovchenko et al. (2021); Ferreira et al. (2025). Within the same finite-temperature NJL framework, it would be highly interesting to investigate the influence of diquark condensates on the evolution of protoquark stars. The corresponding results would be crucial for understanding the cooling mechanism of the compact stars, and could shed light on the behavior of dense matter and its broader cosmological implications.
@@ Acknowledgements.
We thank Prof. Chen Zhang, Prof. Pengcheng Chu, Dr. Bingjun Zuo, and the PKU Pulsar Group for their useful discussions. We also would like to thank the referee for the insightful and valuable comments and suggestions. This work is supported by the National Natural Science Foundation of China (Nos. 12003047, 12133003), the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDB0550300). Wen-Li Yuan is supported by the China Postdoctoral Science Foundation (Grant. No. 2025M773418). Nobutoshi Yasutake is supported by the JSPS KAKENHI (Grant No. 24K07054).